# Cage: Hardware-Accelerated Safe WebAssembly

### v1

@@ 2.1. WebAssembly
To this end, several memory safety solutions have been proposed, which can be divided into three major categories: (i) trip-wire-based approaches (Serebryany et al., 2012, 2023; Bozdoğan et al., 2022; Nethercote and Seward, 2007; Hasabnis et al., 2012; Dang et al., 2015) that employ guard zones around memory objects and allocate designated memory regions, namely shadow memory, that determine whether a memory location is accessible or not, (ii) object-based approaches (Akritidis et al., 2009; Duck and Yap, 2016; Dhurjati et al., 2006; Duck et al., 2017) that track memory safety metadata on a per-object level and ensure memory safety for pointers with respect to an object, and (iii) pointer-based approaches (Nagarakatte et al., 2009; Necula et al., 2002; Jim et al., 2002; Kroes et al., 2018; Corporation, 2013; Woodruff et al., 2014; Kuvaiskii et al., 2017) that keep track of object bounds by either storing them in the pointer itself (e.g., via fat pointers) or in external data structures. Additionally, specialized solutions (Abadi et al., 2009; Kuznetzov et al., 2018; Liljestrand et al., 2019) have also been proposed to mitigate control-flow attacks that occur via memory safety bugs and either analyze applications to enforce valid control flow paths or enforce memory safety for code pointers using hardware- or software-based techniques.
@@ 2.3. Memory Safety Hardware Extensions
As an alternative to flexible yet slow software-based memory safety solutions, CPU designers develop hardware extensions (ARM Ltd., 2019; Qualcomm Technologies, Inc., 2017; Corporation, 2013) to serve as a foundation for efficient memory safety. They provide security primitives to fortify applications with memory safety properties while having minimal memory and performance overheads, making them suitable for production deployment.
@@ Acknowledgement
We thank the anonymous reviewers for their helpful feedback. We also thank Fritz Rehde, Janne Mantyla, and Carlos Chinea Perez for their work and feedback in the early stages of the project. This work was partially supported through a research grant from Huawei Research Finland, the TUM Innovation Network Resilient, Trustworthy, Sustainable (ReTruSt), and an ERC starting grant (ID: 101077577).
@@ A.2. Artifact Check-List (Meta-Information)
Archived (provide DOI)?: 10.5281/zenodo.13347823
@@ A.3.1. How delivered
All source code can be found at the git repositories below, as well as the following persisted DOI for the artifact: 10.5281/zenodo.13347823, which contains all source code as well as the scripts to build and evaluate all artifacts.

### v2

@@ 2.1. WebAssembly
To this end, several memory safety solutions have been proposed, which can be divided into three major categories: (i) trip-wire-based approaches (Serebryany et al., 2012, 2023; Bozdoğan et al., 2022; Nethercote and Seward, 2007; Hasabnis et al., 2012; Dang et al., 2015) that employ guard zones around memory objects and allocate designated memory regions, namely shadow memory, that determine whether a memory location is accessible or not, (ii) object-based approaches (Akritidis et al., 2009; Duck and Yap, 2016; Dhurjati et al., 2006; Duck et al., 2017) that track memory safety metadata on a per-object level and ensure memory safety for pointers with respect to an object, and (iii) pointer-based approaches (Nagarakatte et al., 2009; Necula et al., 2002; Jim et al., 2002; Kroes et al., 2018; Corporation, 2013; Oleksenko et al., 2017; Woodruff et al., 2014; Kuvaiskii et al., 2017) that keep track of object bounds by either storing them in the pointer itself (e.g., via fat pointers) or in external data structures. Additionally, specialized solutions (Abadi et al., 2009; Kuznetzov et al., 2018; Liljestrand et al., 2019) have also been proposed to mitigate control-flow attacks that occur via memory safety bugs and either analyze applications to enforce valid control flow paths or enforce memory safety for code pointers using hardware- or software-based techniques.
@@ 2.3. Memory Safety Hardware Extensions
As an alternative to flexible yet slow software-based memory safety solutions, CPU designers develop hardware extensions (ARM Ltd., 2019; Qualcomm Technologies, Inc., 2017; Corporation, 2013; Oleksenko et al., 2017) to serve as a foundation for efficient memory safety. They provide security primitives to fortify applications with memory safety properties while having minimal memory and performance overheads, making them suitable for production deployment.
@@ Acknowledgement
We thank the anonymous reviewers for their helpful feedback. We also thank Fritz Rehde, Janne Mantyla, and Carlos Chinea Perez for their work and feedback in the early stages of the project. This work was partially supported by a research grant from Huawei Research Finland, the TUM Innovation Network Resilient, Trustworthy, Sustainable (ReTruSt), a Google Safe Compilation grant, and an ERC Starting Grant (ID: 101077577).
@@ A.2. Artifact Check-List (Meta-Information)
Archived (provide DOI)?: 10.5281/zenodo.13772996
@@ A.3.1. How delivered
All source code can be found at the git repositories below, as well as the following persisted DOI for the artifact: 10.5281/zenodo.13772996, which contains all source code as well as the scripts to build and evaluate all artifacts.